OpenStudy (anonymous):

find set values of x for the following inequality |4div(x-1)|>3-(3divx)

OpenStudy (anonymous):

\[|4\div(x-1)|>3-(3\div x)\]

OpenStudy (anonymous):

someone help...

OpenStudy (anonymous):

hang-on... i need to go out for a while.... i will be back in half an hour..... bye...

OpenStudy (anonymous):

Ok, I have worked on this for a while and are hitting many pitfalls. The idea is to split it into two cases, one where x-1 is positive and the absolute signs vanish and one where it is negative and the inequality reverses

OpenStudy (dumbcow):

\[3 - \frac{3}{x} = \frac{3x-3}{x}\] 2 cases \[\frac{4}{x-1} > \frac{3x-3}{x}\] and \[\frac{4}{x-1} < \frac{-(3x-3)}{x}\] Solve each case separately I will look at the negative case first, cross-multiply \[4x <-(3x-3)(x-1)\] \[\rightarrow 3x^{2}-2x+3 < 0\] You will find this has complex solutions so there are no x-values such that its less than 0 next case cross-multiply \[4x > (3x-3)(x-1)\] \[\rightarrow 3x^{2}-10x+3 < 0\] Factor \[(3x-1)(x-3) < 0\] For this product to be negative, one term must be positive, the other negative. This occurs when x is between 1/3 and 3. However, when x=1 the original equation becomes undefined Solution\[x \in (1/3,1) U (1,3)\]

OpenStudy (dumbcow):

i skipped some of the algebra steps do you see how i split the absolute value into 2 cases

OpenStudy (anonymous):

ok... thx... i think i gonna to switch subject... coz tmr i hv chemistry exam... do u know someone good in chemistry?? coz now just 1 online... that's me...

OpenStudy (dumbcow):

shoot, no i don't sorry. i took a couple years of it but i dunno if i can help you

OpenStudy (anonymous):

ok.. nvr mind... i try to review my book.... no notes... ^^

OpenStudy (dumbcow):

good luck